OpenStudy (anonymous):

Choose the letter of the equation which you can use to solve the following word problem. Jill sold half of her comic books and then bought sixteen more. She now has 36. With how many did she begin? Question 3 options: a) 12x+16=36 b) 36+12x=16 c) x+16=36 d) 32x−16=36

OpenStudy (anonymous):

I'm pretty sure the answer is a but I wanna make sure

OpenStudy (anonymous):

Yeah none of these are the right answer The actual equation is\[\frac{ 1 }{ 2 }x+16=36\]

OpenStudy (boldjon):

total number =36 beginning number=n n/2 (sold half)+16 (bought 16 more)=36 (total number) n/2+16-16=36-16 n/2=20 n/2*2=20*2 n=40 the number she began with

rebeccaxhawaii (rebeccaxhawaii):

it would be C because x is what she had originally

OpenStudy (anonymous):

No it wouldn't be @rebeccaxhawaii because 20 is NOT what she had intially It HAS to be 1/2x

rebeccaxhawaii (rebeccaxhawaii):

but what they are looking for is how much she first started with .. how much did she "begin with"

rebeccaxhawaii (rebeccaxhawaii):

that is why the 1/2 equation is not there

OpenStudy (anonymous):

She began with the original number Half of 20 plus 16 is not 36

OpenStudy (anonymous):

You can't solve this equation directly, You could do it like this but you would need a second equation stating that another variable, y, equals 2x Or \[y=2x\]

OpenStudy (anonymous):

And that is not provided

rebeccaxhawaii (rebeccaxhawaii):

yes @Brill is right i misread the question

OpenStudy (anonymous):

thx